OpenStudy (anonymous):

Find the solution of : \(\color{blue}{\frac{\omega}{2} = \cot(114.6 \omega)}\)

OpenStudy (anonymous):

:) It showed up on the mathematics board that it was your first question :P

OpenStudy (anonymous):

Any method except Newton-Raphson and Numerical Methods all those things??

OpenStudy (anonymous):

Really?? Is it my first question? :P

OpenStudy (anonymous):

Apparently c:

OpenStudy (anonymous):

I have asked a hell of questions here.. :)

OpenStudy (anonymous):

its probably your first question in the math subject

OpenStudy (anonymous):

Probably? You are doubting the mathematician techniques of OS.. :P

OpenStudy (anonymous):

i dont understand lol

OpenStudy (anonymous):

so anyways you wanted to solve for w right?

OpenStudy (anonymous):

You were thinking to solve for cot?

OpenStudy (anonymous):

totally

OpenStudy (michele_laino):

we can write: \[\Large 114.6\omega = \operatorname{arccot} \left( {\frac{\omega }{2}} \right)\]

OpenStudy (amoodarya):

'Michele_Laino' you write it as arccot(w/2)=114.6w note that your equation has 1 root but the original question has many roots

OpenStudy (amoodarya):

there is not an interval ? to find roots

OpenStudy (anonymous):

Find any value in any interval.. I just want one value and there is on interval defined..

OpenStudy (anonymous):

*no in place of on..